OpenStudy (anonymous):

Convert the fraction to a decimal. A. 11.01875 B. 11.0316 C. 11.1875 D. 11.316

OpenStudy (anonymous):

And where is the fraction?

OpenStudy (anonymous):

i think d

OpenStudy (anonymous):

\[11\dfrac{3}{16} = 11 + \dfrac{3}{16}\] Now, what is \(\dfrac{3}{16}\)? You know what \(\dfrac{1}{16}\) is, right? multiply it by 3

OpenStudy (anonymous):

3/16

OpenStudy (anonymous):

the answer is 11.1875